Eddy, age 20, plans to deposit $30,000 at age 24.

Raquel plans to deposit $45,000 starting at age 31.

Both expect to earn 7.5 percent per year.

Which plan is better given that neither of these individuals wishes to spend any of their savings until they reach age 60?
Answer #1

Simple interest of Eddy till age of 60 years = $30000 * 37 = $ 1010000

Simple interest of Raquel till age of 60 years = 45000 * 30 = $ 1350000

Hence Raquel's plan is better as he is achieving more savings.

Note :Since both are earning same interest rate of 7.5% per year, we are ignoring the same while computation